What mass of HCl forms from 2.8 g of Cl2?
Step 2: Mass to moles
Step 3: Mole to Mole (Ratio)
Step 4: Mole to Mass
1
Expert's answer
2020-12-06T10:34:29-0500

H2 + Cl2 = 2HCl

M (Cl2) = 70.9 g/mol

M (HCl) = 36.5 g/mol

n (Cl2) = 2.8/70.9 = 0.04 mol

n (HCl) = 2 x n (Cl2) = 2 x 0.04 = 0.08 mol

m (HCl) = 0.08 x 36.5 = 2.9 g